OpenStudy (anonymous):

First, here is the existing map of current structures. It is important that the rollercoaster does not go through the foundation of any of these structures. 1st point: ___6___ 2nd point:___-2___ 3rd point: ___-7___ Using the points above as zeros, construct the polynomial function, f(x), that will be the path of your rollercoaster. Show all of your work. Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences. Solve for the y–intercept for your function, f(x), and then construct a rough graph of your rollercoaster. If your y–intercept is off the graph, give the coordinates of the y–intercept im on number 2 have number 1 answered

OpenStudy (anonymous):

What your question

OpenStudy (anonymous):

First, here is the existing map of current structures. It is important that the rollercoaster does not go through the foundation of any of these structures. 1st point: ___6___ 2nd point:___-2___ 3rd point: ___-7___ Using the points above as zeros, construct the polynomial function, f(x), that will be the path of your rollercoaster. Show all of your work. Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences. Solve for the y–intercept for your function, f(x), and then construct a rough graph of your rollercoaster. If your y–intercept is off the graph, give the coordinates of the y–intercept.

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@Lyssa123 do you see the questions

OpenStudy (anonymous):

yea an I do flvs too

OpenStudy (anonymous):

did you do this assingment and what are you taking?

OpenStudy (anonymous):

math

OpenStudy (anonymous):

algebra 2?

OpenStudy (anonymous):

@ranga help please

OpenStudy (anonymous):

Looks as though neither you nor your graph ever crashes through the roller coaster foundation. However, you do touch the foundation at 3 different x-values: {-7,-2,6}. Hope you're well padded to avoid injury. Those x-values are called "roots" of the polynomial in question. The root -7 becomes the factor (x+7); -2 => (x+2); and 6=> (x-6). Multiply any two of these factors together and simplify the result. Then mult. that result by the remaining factor; the result will the the polynomial you wanted!

OpenStudy (anonymous):

so my equation is f(x)=(x+7)(x+2)(x-6)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now i multiply

OpenStudy (anonymous):

x=6⟹x−6=0⟹(x−6)=0x=−2⟹x+2=0⟹(x+2)=0x=−7⟹x+7=0⟹(x+7)=0(x−6)(x+2)(x+7)=0⟹⟹(x−6)(x+2)(x+7)=original polynomial

OpenStudy (anonymous):

ok im lost lol

OpenStudy (anonymous):

(6,0) and (-2,0) and (-7,0) are points on the x-axis; they're not factors. what I was trying to tell you was that if one root is 6, one factor is (x-6); if another root is -2, the corresponding factor is (x-[-2]). What is the third factor? To get started finding your polynomial, please multiply together (x-6) and (x+2).

OpenStudy (anonymous):

(x-6)(x+2)=x^2-4x-12? is this correct

OpenStudy (anonymous):

so how do i answer the first question

OpenStudy (anonymous):

to finf the function thats what we're doing now

OpenStudy (anonymous):

Right. So, now you have three horiz. intercepts, (6,0) and (-2,0) and (-7,0) , and you have three factors, (x-6),(x+2) and (x+7). Now, with x^2-4x-12 in hand, multiply this polynomial by the third factor, (x+7). Combine like terms. Write your function so that all powers of x are in descending order. Then you'll have your polynomial.

OpenStudy (anonymous):

hold on a sec let me go get a calculator back in 2 min promise

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

back going to do the math now

OpenStudy (anonymous):

ok Your function / your polynomial is f(x)=x^3 + x^2 -28x -12x - 84

OpenStudy (anonymous):

ok and my calculator is acting stupid and let me log this real quick

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so for the first question i log what i multipled right just asking

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so #2 now

OpenStudy (anonymous):

Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences#2

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

@ranga help please @math92130 help

OpenStudy (anonymous):

@m&msdoodle @Brainybeauty

OpenStudy (anonymous):

@camerondoherty @just_one_last_goodbye

OpenStudy (anonymous):

@Luigi0210 i heard you were great can you help me please ill give you medal

OpenStudy (anonymous):

can you help please its urgent

OpenStudy (luigi0210):

@tHe_FiZiCx99

OpenStudy (anonymous):

thanks @luigi0210

OpenStudy (anonymous):

what is the answer to the quation 2. Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences.